Question 7 5 pts XYZ Company purchased a new machine on Ja nuary 1, 2023 for $149,800. The machine had a 7-year life and a $5,600 sa lvage value assigned to it. The machine is being depreciated using the su m-of -the-years -digits method XYZ Company sold the machine on December 31, 2025 for $20,000. Calculate the amount of the loss on the s ale of the machine recorded by XYZ Company
Answer #1
Accumulated depreciation till December 31,2025 = (149800-5600)/28*18= $92700
Loss on the sale of machine = (149800-92700)-20000= $37100